# ROVER: A Multi-Season Dataset for Visual SLAM

### v1

@@ V-C Experiments and Results
We present the results of the experimental evaluation of the SLAM algorithms across a range of sensor configurations and environmental conditions. Each SLAM algorithm was tested using all supported sensor configurations in Table V, combining different cameras and IMUs on all of the 39 sequences recorded. For each SLAM configuration, we compute the ATE for every location, averaged across the four different seasons, along with the overall mean ATE across all locations, the standard deviation of the ATE, and the success rate. To highlight the performance of each SLAM method, we select a representative sensor configuration for each algorithm, specifically the configuration that achieves the lowest mean ATE. For these selected configurations, we further analyze their performance under different environmental conditions, i.e., varying seasons and lighting scenarios, to examine how these factors influence SLAM accuracy.
TABLE VI: Mean Absolute Trajectory Error (ATE) and Success Rate for mono SLAM Configurations across the locations. Each cell presents the mean ATE (in meters) across the five locations – Park, Campus Small, Campus Large, Garden Small, and Garden Large. The second row in each cell shows the mean ATE across all locations, the standard deviation of the ATE, and the Success Rate in percent, calculated as the ratio of valid trajectories to the total number of trials for each configuration. The best results for each location and the overall mean are shown in bold. For the overall mean, results must be reported for each location to be taken into account.
@@
ORB-SLAM3 7.79 | 6.91 | 13.45 | 4.65 | 7.15 x | x | 9.03 | 2.85 | 0.42 2.69 | 1.96 | 11.41 | 0.63 | 2.71
7.82 | 3.86 | 78.95 3.6 | 3.88 | 26.32 4.05 | 4.59 | 94.74
SVO Pro 7.91 | 5.98 | 14.10 | 4.92 | 7.12 7.49 | 6.4 | 13.52 | 4.83 | 7.33 7.58 | 5.84 | 9.27 | 4.72 | 6.68
7.31 | 3.48 | 84.21 7.94 | 3.71 | 100.00 6.78 | 2.75 | 94.74
DROID-SLAM 7.55 | 6.80 | 13.56 | 4.95 | 7.35 5.63 | 5.00 | 9.16 | 3.73 | 5.53 83.98 | 96.36 | 35.59 | x | 86.50
7.76 | 3.46 | 94.74 5.63 | 2.78 | 94.74 77.28 | 45.47 | 26.32
DPVO 7.58 | 6.82 | 13.72 | 4.94 | 7.34 5.36 | 5.41 | 11.86 | 3.38 | 6.04 7.65 | 6.85 | 14.18 | 4.98 | 7.28
7.79 | 3.52 | 94.74 6.47 | 3.48 | 100.00 8.22 | 3.9 | 100.00
DPV-SLAM 7.52 | 6.86 | 13.66 | 4.88 | 7.31 5.75 | 5.25 | 11.29 | 3.14 | 5.66 7.58 | 6.87 | 14.22 | 4.99 | 7.27
7.77 | 3.51 | 94.74 6.24 | 3.39 | 100.00 8.21 | 3.91 | 100.00
TABLE VII: Mono-inertial SLAM results. An ’x’ indicates that no valid trajectories were generated by the SLAM configuration.
@@
OpenVINS 1.10 | 1.92 | 3.38 | 0.94 | 1.85 20.46 | 1.12 | 2.45 | 0.47 | 0.83 x | 89.24 | 2.16 | x | 0.56 1.11 | 1.23 | 3.67 | 0.75 | 1.13 1.04 | 1.52 | 3.92 | 0.86 | 1.55
1.74 | 1.13 | 78.95 4.14 | 10.94 | 68.42 23.53 | 45.88 | 21.05 1.49 | 1.13 | 94.74 1.82 | 1.30 | 100.00
VINS-Fusion x | x | x | x | x 5.65 | 4.08 | 6.67 | 2.85 | 3.50 x | x | x | x | x x | x | x | 17.75 | 4.96 x | x | 6.92 | 3.42 | 2.29
x | x | 0.00 4.14 | 1.75 | 73.68 x | x | 0.00 8.16 | 8.02 | 21.05 4.89 | 4.69 | 21.05
ORB-SLAM3 x | x | x | x | x x | x | x | x | x x | 13.42 | x | x | 17.21 6.70 | 4.27 | x | 52.71 | x x | 3.45 | x | 2.55 | 2.19
x | x | 0.00 x | x | 0.00 15.32 | 15.55 | 10.53 19.15 | 34.96 | 36.84 2.69 | 1.77 | 21.05
SVO Pro 1.54 | 1.95 | 11.97 | 1.90 | 3.71 x | x | x | x | x 7.11 | 32.72 | x | x | x 1.80 | 3.72 | 4.98 | 1.06 | 2.62 x | 5.34 | 93.31 | 1.35 | 5.98
3.29 | 3.46 | 57.89 x | x | 0.00 24.19 | 25.29 | 15.79 2.62 | 2.42 | 63.16 15.58 | 32.22 | 42.11
@@
TABLE VIII: Stereo and stereo-inertial SLAM results. ’n/a’ indicates that the sensor configuration is not supported by the SLAM.
SLAM Stereo Stereo-inertial
T265 T265 - internal T265 - external
OpenVINS n/a 1.11 | 1.54 | 3.32 | 0.77 | 0.81 1.29 | 1.11 | 4.10 | 0.73 | 1.36
1.43 | 1.05 | 94.74 1.74 | 1.55 | 100.00
VINS-Fusion x | x | x | x | x 3.95 | 2.86 | 5.58 | 1.63 | 2.34 x | 3.26 | x | x | x
x | x | 0.00 3.13 | 2.09 | 73.68 3.26 | 0.00 | 5.26
ORB-SLAM3 2.78 | 0.82 | 5.57 | 0.98 | 0.79 32.05 | 0.58 | x | 0.39 | 0.59 0.54 | 1.07 | x | 0.31 | 0.67
2.41 | 3.32 | 84.21 7.51 | 21.02 | 47.37 0.71 | 0.73 | 42.11
SVO Pro 57.12 | 41.91 | 99.44 | 32.11 | 53.83 2.62 | 11.37 | 3.35 | 0.97 | 1.43 1.79 | 1.11 | 3.58 | 1.01 | 1.04
49.89 | 27.91 | 73.68 4.21 | 8.24 | 68.42 1.7 | 1.26 | 100.00
DROID-SLAM 83.98 | 96.36 | 35.59 | x | 86.50 n/a n/a
77.28 | 45.47 | 26.32
TABLE IX: RGB-D and RGB-D-inertial SLAM results.
SLAM RGB-D RGB-D-inertial
D435i D435i - internal D435i - external
ORB-SLAM3 1.84 | 3.37 | 0.53 | 0.61 | 0.64 2.84 | 0.54 | 36.52 | 0.48 | 0.34 2.06 | 0.97 | 62.65 | 0.75 | 12.63
1.26 | 2.11 | 94.74 5.94 | 19.28 | 73.68 8.35 | 17.49 | 78.95
DROID-SLAM 1.85 | 1.87 | 3.81 | 2.59 | 1.72 n/a n/a
2.32 | 2.89 | 94.74
@@
The evaluation of SLAM configurations reveals several important trends across different sensor setups and algorithms. In the mono configurations, see Table VI, there is a general trend of higher ATE values across all camera types, with ORB-SLAM3 using the D435i camera achieving the best results. However, this configuration suffers from a low success rate, indicating instability. In contrast, the T265 camera shows relatively good results with a much higher success rate, although it struggles in larger environments. For mono-inertial setups, depicted in Table VII, the results are more varied. Some configurations fail to produce valid results, while others perform relatively well. OpenVINS with the T265 camera in the internal configuration stands out as one of the better-performing setups, showing solid results across different environments. In stereo configurations, see Table VIII, there are significant difficulties across most setups, except for ORB-SLAM3, which achieves better consistency and accuracy. When adding inertial data to stereo configurations, the results become more stable. OpenVINS performs well with both internal and external T265 configurations, and ORB-SLAM3 shows good accuracy, though its success rate remains lower. Overall, stereo-inertial setups provide more reliable performance than pure stereo configurations. The results for the RGB-D configuration are presented in Table IX, demonstrating a significant improvement in performance. ORB-SLAM3 and DROID-SLAM show good accuracy and consistent results across small and large environments alike. The inclusion of depth data appears to enhance SLAM performance considerably, making RGB-D setups among the most robust options tested. However, when inertial data is added to RGB-D configurations, the results do not improve further; in fact, the performance tends to decline. Both the accuracy and success rate drop compared to RGB-D-only configurations, suggesting that the addition of inertial data, in this case, does not contribute positively to the results.
@@
TABLE X: Performance of the best-performing sensor configuration for each SLAM method across different seasons. Each cell shows the performance metrics for the best configuration, with OV representing OpenVINS, VF for VINS-Fusion, ORB for ORB-SLAM3, SVO for SVO Pro, DROID for DROID-SLAM, and DPV for DPV-SLAM. Configurations are denoted as follows: S-I for stereo-inertial, D for RGB-D, M for mono, I for internal IMU, and E for external IMU. Please note that the dataset does not include a winter sequence for the park location.
@@
OV (S-I T265 I) 1.51 | 1.22 | 3.88 | 0.71 | 0.61 1.04 | 1.16 | x | 0.89 | 0.92 n/a | 2.57 | 2.99 | 0.68 | 0.92 0.78 | 1.22 | 3.09 | 0.82 | 0.79 1.11 | 1.54 | 3.32 | 0.77 | 0.81
1.59 | 1.33 | 100.00 1.00 | 0.12 | 80.00 1.79 | 1.55 | 100.00 1.34 | 1.00 | 100.00 1.43 | 1.05 | 94.74
VF (S-I T265 I) 6.62 | x | 6.34 | 1.53 | 2.34 x | 2.46 | x | 2.14 | 2.35 n/a | 3.25 | 4.50 | 1.23 | 2.34 1.28 | x | 5.89 | 1.63 | x 3.95 | 2.86 | 5.58 | 1.63 | 2.34
4.21 | 2.65 | 80.00 2.31 | 0.17 | 60.00 2.83 | 2.15 | 100.00 2.93 | 2.56 | 60.00 3.13 | 2.09 | 73.68
ORB (D D435i) 2.56 | 9.21 | 0.64 | 0.64 | 0.86 2.40 | 0.42 | 0.63 | 0.88 | 0.95 n/a | x | 0.42 | 0.46 | 0.39 0.57 | 0.47 | 0.44 | 0.48 | 0.36 1.84 | 3.37 | 0.53 | 0.61 | 0.64
2.78 | 3.68 | 100.00 1.06 | 0.78 | 100.00 0.42 | 0.30 | 75.00 0.46 | 0.07 | 100 1.26 | 2.11 | 94.74
SVO (S-I T265 E) 1.88 | 1.11 | 5.51 | 0.92 | 1.11 1.08 | 1.08 | 3.15 | 1.12 | 1.06 n/a | 1.15 | 2.28 | 0.80 | 1.04 2.40 | 1.12 | 3.36 | 1.20 | 0.95 1.79 | 1.11 | 3.58 | 1.01 | 1.04
2.11 | 1.94 | 100.00 1.50 | 0.93 | 100.00 1.32 | 1.01 | 100.00 1.81 | 1.04 | 100.00 1.70 | 1.26 | 100.00
DROID (D D435i) 1.95 | 5.23 | x | 3.02 | 5.28 3.16 | 1.58 | 10.22 | 6.73 | 0.38 n/a | 0.32 | 0.53 | 0.22 | 0.29 0.45 | 0.35 | 0.70 | 0.37 | 0.93 1.85 | 1.87 | 3.81 | 2.59 | 1.72
3.87 | 1.66 | 80.00 4.42 | 4.03 | 100.00 0.34 | 0.24 | 100.00 0.56 | 0.25 | 100.00 2.32 | 2.89 | 94.74
DPVO (M D435i) 5.10 | 5.42 | 13.05 | 2.16 | 6.25 6.20 | 3.83 | 12.68 | 3.58 | 6.38 n/a | 6.34 | 10.02 | 3.90 | 5.68 4.78 | 6.06 | 11.67 | 3.88 | 5.86 5.36 | 5.41 | 11.86 | 3.38 | 6.04
6.40 | 4.03 | 100.00 6.53 | 3.67 | 100.00 6.48 | 4.54 | 100.00 6.45 | 3.05 | 100.00 6.47 | 3.48 | 100.00
DPV (M D435i) 5.26 | 5.25 | 12.79 | 2.34 | 5.72 7.10 | 3.57 | 12.64 | 2.67 | 5.98 n/a | 5.85 | 8.37 | 3.50 | 5.53 4.89 | 6.34 | 11.38 | 4.04 | 5.40 5.75 | 5.25 | 11.29 | 3.14 | 5.66
6.27 | 3.88 | 100.00 6.39 | 3.92 | 100.00 5.81 | 3.90 | 100.00 6.41 | 2.90 | 100.00 6.24 | 3.39 | 100.00
Mean 3.55 | 4.57 | 7.03 | 1.62 | 3.17 3.50 | 2.01 | 7.86 | 2.57 | 2.57 n/a | 3.25 | 4.16 | 1.54 | 2.31 2.16 | 2.59 | 5.22 | 1.77 | 2.38 3.09 | 3.06 | 5.71 | 1.88 | 2.61
3.89 | 1.13 | 94.29 3.32 | 1.83 | 91.43 2.71 | 1.70 | 96.43 2.85 | 1.26 | 94.29 3.22 | 0.97 | 93.99
TABLE XI: Performance of the best-performing sensor configuration for each SLAM method across different lighting conditions.
@@
OV (S-I T265 I) 7.11 | 1.57 | 1.15 | 0.72 | 0.78 0.98 | 2.05 | 0.96 | 1.31 | 0.96 1.14 | 0.92 | 3.57 | 0.80 | 1.23 3.65 | 2.56 | 2.21 | 3.00 | 1.20 3.22 | 1.78 | 1.97 | 1.46 | 1.04
2.26 | 2.73 | 100.00 1.25 | 0.47 | 100.00 1.53 | 1.15 | 100.00 2.53 | 0.91 | 100 1.89 | 1.53 | 100.00
VF (S-I T265 I) 1.64 | x | 4.55 | 1.70 | 2.40 x | x | 4.95 | 4.73 | 1.82 x | x | 29.17 | 13.18 | 15.83 x | x | 4.48 | 2.90 | 4.47 1.64 | x | 10.79 | 5.63 | 6.13
2.57 | 1.36 | 80.00 3.83 | 1.75 | 60.00 19.4 | 8.57 | 60.00 3.95 | 0.91 | 60.00 7.06 | 7.95 | 65.00
ORB (D D435i) 0.50 | 0.45 | 0.51 | 0.45 | 0.41 4.88 | 8.30 | 5.98 | 5.40 | 6.28 6.85 | 8.22 | 10.97 | x | x 5.85 | 8.84 | 3.79 | x | 0.66 4.52 | 6.45 | 5.31 | 2.93 | 2.45
0.47 | 0.04 | 100.00 6.17 | 1.31 | 100.00 8.68 | 2.10 | 60.00 4.78 | 3.44 | 80.00 4.61 | 3.53 | 85.00
SVO (S-I T265 E) 1.85 | 1.02 | 2.66 | 1 .00| 1.15 1.95 | 1.12 | 3.61 | 1.18 | 0.99 3.89 | 1.21 | 3.90 | x | 1.61 27.96 | 1.09 | x | 1.22 | 2.24 8.91 | 1.11 | 3.39 | 1.13 | 1.50
1.54 | 0.72 | 100.00 1.77 | 1.10 | 100.00 2.65 | 1.44 | 80.00 8.13 | 13.23 | 80.00 3.31 | 6.23 | 90.00
DROID (D D435i) 0.36 | 0.24 | 0.44 | 0.25 | 0.29 0.58 | 0.34 | 0.47 | 1.42 | 3.58 5.61 | 1.32 | 2.68 | 8.11 | 10.40 0.29 | 0.57 | 0.29 | 0.60 | 0.78 1.71 | 0.62 | 0.97 | 2.60 | 3.76
0.32 | 0.08 | 100.00 1.28 | 1.35 | 100.00 5.62 | 3.75 | 100.00 0.51 | 0.21 | 100.00 1.93 | 2.88 | 100.00
DPVO (M D435i) 2.67 | 6.40 | 8.76 | 4.47 | 5.15 4.87 | 6.11 | 8.92 | 4.40 | 5.22 7.21 | 5.91 | 10.78 | 4.78 | 6.56 2.43 | 5.59 | 9.25 | 4.25 | 5.22 4.29 | 6.00 | 9.43 | 4.48 | 5.54
5.49 | 2.27 | 100.00 5.90 | 1.80 | 100.00 7.05 | 2.27 | 100.00 5.35 | 2.50 | 100.00 5.95 | 2.15 | 100.00
DPV (M D435i) 3.46 | 5.80 | 8.98 | 4.11 | 4.99 4.66 | 5.39 | 8.79 | 4.46 | 4.81 7.54 | 5.42 | 10.82 | 5.31 | 6.65 2.80 | 4.52 | 8.05 | 3.82 | 4.24 4.62 | 5.29 | 9.16 | 4.43 | 5.17
5.47 | 2.15 | 100.00 5.62 | 1.80 | 100.00 7.15 | 2.25 | 100.00 4.69 | 1.99 | 100.00 5.73 | 2.10 | 100.00
Mean 2.51 | 2.58 | 3.86 | 1.81 | 2.17 2.99 | 3.88 | 4.81 | 3.27 | 3.38 5.37 | 3.83 | 10.27 | 6.44 | 7.05 7.16 | 3.86 | 4.68 | 2.63 | 2.69 4.13 | 3.54 | 5.86 | 3.24 | 3.66
2.59 | 1.09 | 97.14 3.69 | 0.48 | 94.29 7.44 | 2.56 | 85.71 4.28 | 4.51 | 88.57 4.35 | 2.41 | 91.43
In addition to the overall performance evaluation of the SLAM algorithms, we also examine the impact of environmental factors, including seasons and lighting conditions, on the best-performing configurations. Seasonal changes play a critical role in the performance of SLAM systems, as shown in Table X. Winter and spring generally offer the most favorable conditions for accurate localization, with lower vegetation density and more stable lighting reducing visual complexity. In contrast, summer and autumn, characterized by denser vegetation and more variable lighting, lead to significantly higher ATE values. Systems such as ORB-SLAM3 and DROID-SLAM perform well during winter and spring, but encounter increased ATEs in summer, reflecting the seasons’ characteristics. Notably, OpenVINS (stereo-inertial T265 - internal) demonstrates strong year-round consistency, while ORB-SLAM3 and DROID-SLAM (RGB-D D435i) excel in winter and spring. Lighting conditions also have a profound impact on SLAM performance.
The results show that SLAM systems perform most reliably in daylight, where they achieve the lowest ATE and the highest success rates, as shown in Table XI. For example, DROID-SLAM (RGB-D D435i) consistently performs well during daytime across various locations, while OpenVINS (stereo-inertial T265 internal) also exhibits stable performance in both daylight and dusk conditions. As light decreases in dusk and night scenarios, performance declines sharply for most algorithms, with night conditions posing the greatest challenge—though OpenVINS shows relative resilience. Many SLAM methods face difficulties maintaining accuracy in low-light environments, as evidenced by higher error rates and lower success rates at night. ORB-SLAM3 (RGB-D D435i), for instance, shows a sharp increase in ATE during night conditions despite performing well during the day. Even when external lighting is introduced, as in the night + light scenario, the improvement is only marginal, highlighting the need for enhanced robustness under low-visibility conditions. These results indicate that while some configurations, like DROID-SLAM and OpenVINS, are more resilient under suboptimal lighting, many SLAM methods still struggle with low-light feature extraction, especially in larger environments.

### v2

@@ V-C Experiments and Results
We present the results of the experimental evaluation of the SLAM algorithms across a range of sensor configurations and environmental conditions. Each SLAM algorithm was tested using all supported sensor configurations in Table VI, combining different cameras and IMUs on all of the 39 sequences recorded. For each SLAM configuration, we compute the ATE and RPE for every location (averaged across the four seasons) and then determine the overall mean ATE (mATE), mean RPE (mRPE), and SR across all locations. To highlight the performance of each SLAM method, we select a representative sensor configuration for each algorithm, specifically the configuration that achieves the lowest mATE. For these selected configurations, we further analyze their performance under different environmental conditions, i.e., varying seasons and lighting scenarios, to examine how these factors influence SLAM accuracy.
TABLE VII: mATE [m], mRPE [m] and SR [%] for mono SLAM configurations. The best results for mATE and mRPE are shown in bold.
@@
mATE mRPE SR mATE mRPE SR mATE mRPE SR
ORB-SLAM3 7.82 4.05 78.95 3.60 1.34 26.32 4.05 0.50 94.74
SVO Pro 7.31 2.94 84.21 7.94 3.29 100.00 6.78 5.56 94.74
DROID-SLAM 7.76 4,24 94.74 5.63 1,73 94.74 77.28 4.75 26.32
DPVO 7.79 3.43 94.74 6.47 2,15 100.00 8.22 5.48 100.00
DPV-SLAM 7.77 1.63 94.74 6.24 1.84 100.00 8.21 5.70 100.00
TABLE VIII: Mono-inertial SLAM results. An ’x’ indicates that no valid trajectories were generated by the SLAM configuration.
@@
mATE mRPE SR mATE mRPE SR mATE mRPE SR mATE mRPE SR mATE mRPE SR
OpenVINS 1.74 0.17 78.95 4.14 0.19 68.42 23.53 0.24 21.05 1.49 0.16 94.74 1.82 0.15 100.00
VINS-Fusion x x 0.00 4.14 0.98 73.68 x x 0.00 8.16 0.83 21.05 4.89 0.32 21.05
ORB-SLAM3 x x 0.00 x x 0.00 15.32 1.54 10.53 19.15 1.46 36.84 2.69 0.38 21.05
SVO Pro 3.29 0.18 57.89 x x 0.00 24.19 0.26 15.79 2.62 0.17 63.16 15.58 0.19 42.11
@@
TABLE IX: Stereo and stereo-inertial SLAM results. ’n/a’ indicates that the sensor configuration is not supported by the SLAM.
SLAM Stereo Stereo-inertial (internal) Stereo-inertial (external)
mATE mRPE SR mATE mRPE SR mATE mRPE SR
OpenVINS n/a n/a n/a 1.43 0.23 94.74 1.74 0.27 100.00
VINS-Fusion x x 0.00 3.13 0.57 73.68 3.26 0.28 5.26
ORB-SLAM3 2.41 0.35 84.21 7.51 0.48 47.37 0.71 0.14 42.11
SVO Pro 49.89 0.38 73.68 4.21 0.17 68.42 1.70 0.16 100.00
DROID-SLAM 77.28 0.76 26.32 n/a n/a n/a n/a n/a n/a
TABLE X: RGBD and RGBD-inertial SLAM results.
SLAM RGBD RGBD-inertial (internal) RGBD-inertial (external)
mATE mRPE SR mATE mRPE SR mATE mRPE SR
ORB-SLAM3 1.26 0.17 94.74 5.94 0.18 73.68 8.35 0.33 78.95
DROID-SLAM 2.32 0.21 94.74 n/a n/a n/a n/a n/a n/a
@@
The evaluation of SLAM configurations reveals several important trends across different sensor setups and algorithms. In the mono configurations (see Table VII), there is a general trend of higher mATE and mRPE values across all camera types. Notably, ORB-SLAM3 using the D435i camera achieves the lowest mATE, which indicates the best absolute pose estimation. However, this configuration suffers from a low SR and poor mRPE, suggesting instability. In contrast, the T265 camera shows better mRPE and a much higher SR, although it exhibits a higher mATE. For mono-inertial setups, depicted in Table VIII, the results are more varied. Some configurations fail to produce valid results, while others perform relatively well. OpenVINS with the T265 camera in the internal configuration stands out as one of the better-performing setups, showing solid results in both mATE and mRPE. In stereo configurations (see Table IX), most setups experience significant difficulties in achieving low mATE values, while the mRPE results remain largely comparable. ORB-SLAM3 stands out by attaining better consistency and higher overall accuracy in both mATE and mRPE. When adding inertial data to stereo configurations the results become more stable. Both OpenVINS, which performs well with both internal and external T265 configurations, and SVO Pro using an external IMU exhibit comparable performance. ORB-SLAM3 with the external IMU stands out as the best performing method by achieving the lowest mATE and mRPE, but its relatively low SR makes it a rather unstable method. Overall, stereo-inertial setups offer more reliable localization performance than pure stereo configurations. The results for the RGBD configuration are presented in Table X, demonstrating a significant improvement in performance. The inclusion of depth data enhances SLAM performance considerably, with RGBD configurations showing good localization results in terms of both mATE and mRPE, and ORB-SLAM3 achieving a high SR, which makes it one of the most robust options tested. However, when inertial data is added to RGBD configurations, the results do not improve further; in fact, the performance tends to decline. Both the accuracy and SR drop compared to RGBD-only configurations, suggesting that the addition of inertial data, in this case, does not contribute positively to the results.
@@
TABLE XI: Performance of the best-performing sensor configuration for each SLAM method across different seasons, with OV representing OpenVINS, VF for VINS-Fusion, ORB for ORB-SLAM3, SVO for SVO Pro, DROID for DROID-SLAM, and DPV for DPV-SLAM. Sensor configurations are denoted as follows: S-I for stereo-inertial, D for RGBD, M for mono, I for internal IMU, and E for external IMU.
@@
mATE mRPE SR mATE mRPE SR mATE mRPE SR mATE mRPE SR mATE mRPE SR
OV (S-I T265 I) 1.59 0.17 100.00 1.00 0.33 80.00 1.79 0.28 100.00 1.34 0.17 100.00 1.43 0.23 94.74
VF (S-I T265 I) 4.21 0.82 80.00 2.31 0.40 60.00 2.83 0.56 100.00 2.93 0.40 60.00 3.13 0.57 73.68
ORB (D D435i) 2.78 0.19 100.00 1.06 0.17 100.00 0.42 0.16 75.00 0.46 0.17 100.00 1.26 0.17 94.74
SVO (S-I T265 E) 2.11 0.18 100.00 1.50 0.16 100.00 1.32 0.14 100.00 1.81 0.17 100.00 1.70 0.16 100.00
DROID (D D435i) 3.87 0.43 80.00 4.42 0.24 100.00 0.34 0.07 100.00 0.56 0.08 100.00 2.32 0.21 94.74
DPVO (M D435i) 6.40 2.13 100.00 6.53 2.04 100.00 6.48 2.02 100.00 6.45 2.39 100.00 6.47 2.15 100.00
DPV (M D435i) 6.27 1.67 100.00 6.39 1.93 100.00 5.81 1.87 100.00 6.41 1.88 100.00 6.24 1.84 100.00
Mean 3.89 0.80 94.29 3.32 0.75 91.43 2.71 0.73 96.43 2.85 0.75 94.29 3.22 0.76 93.99
TABLE XII: Performance of the best-performing sensor configuration for each SLAM method across different lighting conditions.
@@
mATE mRPE SR mATE mRPE SR mATE mRPE SR mATE mRPE SR mATE mRPE SR
OV (S-I T265 I) 2.26 0.17 100.00 1.25 0.15 100.00 1.53 0.19 100.00 2.53 0.40 100.00 1.89 0.23 100.00
VF (S-I T265 I) 2.57 0.50 80.00 3.83 0.54 60.00 19.40 0.75 60.00 3.95 0.98 60.00 7.06 0.68 65.00
ORB (D D435i) 0.47 0.17 100.00 6.17 0.85 100.00 8.68 0.17 60.00 4.78 0.15 80.00 4.61 0.36 85.00
SVO (S-I T265 E) 1.54 0.16 100.00 1.77 0.19 100.00 2.65 0.21 80.00 8.13 0.10 80.00 3.31 0.17 90.00
DROID (D D435i) 0.32 0.07 100.00 1.28 0.18 100.00 5.62 0.15 100.00 0.51 0.11 100.00 1.93 0.13 100.00
DPVO (M D435i) 5.49 2.09 100.00 5.90 2.88 100.00 7.05 4.11 100.00 5.35 1.98 100.00 5.95 2.77 100.00
DPV (M D435i) 5.47 1.80 100.00 5.62 2.75 100.00 7.15 3.33 100.00 4.69 1.51 100.00 5.73 2.35 100.00
Mean 2.59 0.71 97.14 3.69 1.08 94.29 7.44 1.27 85.71 4.28 0.76 88.57 4.35 0.96 91.43
Figure 8: Comparison of best-performing SLAM methods regarding mATE and mRPE under different lighting conditions and seasons.
@@
In addition to the overall performance evaluation of the SLAM algorithms, we also examine the impact of environmental factors, including seasons and lighting conditions, on the best-performing configurations. Seasonal changes play a critical role in the performance of SLAM systems, as shown in Table XI and visualized in Figure 8. Winter and spring generally offer the most favorable conditions for accurate localization, with lower vegetation density and more stable lighting reducing visual complexity. In contrast, summer and autumn, characterized by denser vegetation and more variable lighting, lead to significantly higher mATE values. While the absolute errors (mATE) are clearly affected by seasonal variations, the relative errors (mRPE) remains only minimally impacted. Systems such as ORB-SLAM3 and DROID-SLAM perform well during winter and spring, but encounter increased mATEs in summer, reflecting the seasons’ characteristics. Notably, OpenVINS (stereo-inertial T265 - internal) demonstrates strong year-round consistency, while ORB-SLAM3 and DROID-SLAM (RGBD D435i) excel in winter and spring. Lighting conditions also have a profound impact on SLAM performance.
The results show that SLAM systems perform most reliably in daylight, where they achieve the lowest mATE and the highest SR, as shown in Table XII. For example, DROID-SLAM (RGBD D435i) consistently performs well during daytime across various locations, while OpenVINS (stereo-inertial T265 internal) also exhibits stable performance in both daylight and dusk conditions. As light decreases in dusk and night scenarios, both mATE and mRPE performance declines sharply for most algorithms, with night conditions posing the greatest challenge – though OpenVINS shows relative resilience. Many SLAM methods face difficulties maintaining accuracy in low-light environments, as evidenced by higher error rates and lower SR at night. ORB-SLAM3 (RGBD D435i), for instance, shows a sharp increase in ATE during night conditions despite performing well during the day. Even when external lighting is introduced, as in the night + light scenario, the improvement is only marginal, highlighting the need for enhanced robustness under low-visibility conditions. These results indicate that while some configurations, like DROID-SLAM and OpenVINS, are more resilient under suboptimal lighting, many SLAM methods still struggle with low-light feature extraction, especially in larger environments.